# Design Space Exploration of Embedded SoC Architectures for Real-Time Optimal Control

### v1

@@ VI-B Area Evaluation
To understand area synthesis was run on Saturn and Gemmini are both measured in the ASAP7 toolkit [6]. We found that both Gemmini and Saturn consume relatively smaller areas than an out-of-order MegaBOOM core with 2FPUs[29]. Gemmini has smaller areas overhead (1.5-2mm2). A breakdown of the area for a 4$\times$4 output stationary Gemmini with a 32KB scratchpad and a V512D256 reference Saturn, both driven by a Rocket core, is depicted in Figure 21. Proportionally, the FPFMAs in Gemmini’s mesh combined with the scratchpad contribute to a greater proportion of the design than in Saturn. However, this is because Saturn also supports a vectorized integer pipeline. Additionally, Gemmini’s scratchpad has 16x the capacity of the Saturn register file despite only using 35% more area. This is because the scratchpad is synthesized with SRAMs, whereas the Saturn design uses flip-flops. This suggests that Saturn’s area could be significantly improved by using a more dense memory technology for the register file and by replacing the integer backend with a simplified scalar design to better align with the computing requirements of robotics workloads.
@@
Additional area report Table II provides a comparison of Gemmini with and without our GEMV architectural change. The additional 2% area increment in a 4x4 mesh is attributed to the extra scratchpad banks since our feature needs at least $DIM+1$ banks to ensure all the necessary data for the mesh is loaded in one cycle. In all runs, we configured Gemmini to have $DIM+1$ scratchpad banks, rounded up to the nearest power of 2. This is a current requirement of a legal Gemmini configuration; however, we would like to address this constraint in the future. Currently, as $DIM$ scales, the utilization of these extra banks decreases, making a 16x16 mesh where 15 banks are in excess impractical. This can be attributed to the extra logic per scratchpad bank, as traditional Gemmini only utilizes the 2 banks in a single computer. The area is largely dominated by scratchpad banks, which would benefit from the changes stated above. Looking further, the mesh area has been insignificantly impacted by our design due to the lack of logic between the mesh columns. On the other hand, ExecuteController grows 9.2% in size between 4x4 GEMM and 4x4 GEMV as it shoulders more responsibility to distribute data from different scratchpad banks to every PE. This relationship appears to be linear as DIM scales, as the area grows by 18% between 8x8 GEMM and 8x8 GEMV.
@@ VII Conclusion
Our profiling and optimization across various computing platforms ranging from CPUs and vector machines to domain-specialized accelerators encompass both kernel-level benchmarks and a comprehensive end-to-end robotic workload. We rigorously compare integrated accelerators like RISC-V cores with vector extensions to more decoupled systems such as systolic arrays, evaluating their performance, area efficiency, and utilization. This exploration not only quantifies the trade-offs inherent between these architectural paradigms but also underscores how the choice of hardware architecture is contingent upon specific workload characteristics and application demands.
This project conducted a thorough design space exploration for architecture research for classical and optimization-based algorithms for embedded robotics applications. Different architectures are optimal under each area constraint, and well-optimized software mappings are critical to achieve improved performance. To address these challenges and open questions, future work will involve design-space exploration across a broader spectrum of hardware design spaces, as well as robotics algorithms. Furthermore, to facilitate rapid evaluation of new hardware designs, we are currently building upon the lessons learned in this work to develop automated code-generation flows to emit optimized embedded solvers on top of the matlib interface, with the end goal of being able to pass in hardware configurations and robot parameters (which impact matrix and vector sizes), generating optimized libraries for the desired targets.
VIII Appendix
@@
    • https://github.com/kris-dong/matlib
    • https://github.com/ucb-bar/Baremetal-NN/

### v2

@@ VI-B Area Evaluation
To understand area usage, we use post-synthesis area results generated by ASAP7 toolkit [6]. We found that both Gemmini and Saturn consume relatively smaller areas than an out-of-order MegaBOOM core with 2FPUs[27]. Gemmini has smaller areas overhead (1.5-2mm2). A breakdown of the area for a 4$\times$4 output stationary Gemmini with a 32KB scratchpad and a V512D256 reference Saturn, both driven by a Rocket core, is depicted in Figure 21. Proportionally, the FPFMAs in Gemmini’s mesh combined with the scratchpad contribute to a greater proportion of the design than in Saturn. However, this is because Saturn also supports a vectorized integer pipeline. Additionally, Gemmini’s scratchpad has 16$\times$ the capacity of the Saturn register file despite only using 35% more area. This is because the scratchpad is synthesized with SRAMs, whereas the Saturn design uses flip-flops. This suggests that Saturn’s area could be significantly improved by using a more dense memory technology for the register file and by replacing the integer backend with a simplified scalar design to better align with the computing requirements of robotics workloads.
Additional area report Table II provides a comparison of Gemmini with and without our GEMV architectural change. The additional 2% area increment in a 4x4 mesh is attributed to the extra scratchpad banks since our feature needs at least $\textrm{DIM}+1$ banks to ensure all the necessary data for the mesh is loaded in one cycle. In all runs, we configured Gemmini to have $\textrm{DIM}+1$ scratchpad banks, rounded up to the nearest power of 2. This is a current requirement of a legal Gemmini configuration; however, we would like to address this constraint in the future. Currently, as DIM scales, the utilization of these extra banks decreases, making a 16x16 mesh where 15 banks are in excess impractical. This can be attributed to the extra logic per scratchpad bank, as traditional Gemmini only utilizes the 2 banks in a single computer. The area is largely dominated by scratchpad banks, which would benefit from the changes stated above. Looking further, the mesh area has been insignificantly impacted by our design due to the lack of logic between the mesh columns. On the other hand, ExecuteController grows 9.2% in size between 4$\times$4 GEMM and 4$\times$4 GEMV as it shoulders more responsibility to distribute data from different scratchpad banks to every PE. This relationship appears to be linear as DIM scales, as the area grows by 18% between 8$\times$8 GEMM and 8$\times$8 GEMV.
@@ VII Conclusion
Our profiling and optimization across various computing platforms ranging from CPUs and vector machines to domain-specialized accelerators encompass both kernel-level benchmarks and a comprehensive end-to-end robotic workload. We compare integrated accelerators like RISC-V cores with vector extensions to more decoupled systems such as systolic arrays, evaluating their performance, area efficiency, and utilization. This exploration not only quantifies the trade-offs inherent between these architectural paradigms but also underscores how the choice of hardware architecture is contingent upon specific workload characteristics and application demands.
This project conducts a thorough design space exploration for architecture research for classical and optimization-based algorithms for embedded robotics applications. Different architectures are optimal under each area constraint, and well-optimized software mappings are critical to achieve improved performance. To address these challenges and open questions, future work will involve design-space exploration across a broader spectrum of hardware design spaces, as well as robotics algorithms.
A significant challenge for deploying optimized robotics algorithms on specialized hardware is the engineering overhead of hand-optimizing the software mapping. To facilitate rapid evaluation of new hardware designs, we pursue two potential options. First, we explore a library-based approach supporting common operations. Existing interfaces such as Eigen [10] result in poor utilization for small operands, and do not provide clean abstractions for specialized memory units such as software-managed scratchpads.
To address this, we develop the matlib library, with ongoing work on applying optimizations such as kernel fusion. Currently matlib is implemented using RVV intrinsic, and has planned support for the Gemmini ISA. We are currently developing automated code-generation flows to emit optimized embedded solvers on top of the matlib interface, with the end goal of being able to pass in hardware configurations and robot parameters (which impact matrix and vector sizes), generating optimized libraries for the desired targets. A second approach is to express robotic algorithms directly as computation graphs in an intermediate representation such as an MLIR [lattner2021mlir] dialect and adapt end-to-end code-generation techniques.
Additionally, realistic robotics workloads consist of multiple software nodes executing concurrently on an embedded SoC. Given optimized mappings of multiple algorithms to several hardware backends, future work will consider strategies for dynamically scheduling these systems while addressing physical and computational constraints in robotic systems.
APPENDIX
@@
    • https://github.com/ucb-bar/matlib